OpenStudy (anonymous):

Solve: -4(x + 10) - 6 = -3(x - 2) A. -40 B. -46 C. -52 D. 52 @britneyspears

OpenStudy (anonymous):

-4x-40-6=-3x+6 -4x-46=-3x+6 -x=52 x=-52

OpenStudy (anonymous):

\[-4(x+10)-6 = -3(x-2)\] distribute both sides \[-4x -40-6=-3x+6\] add 3x to both sides and add 40 to both sides \[-x = 52\] Divide both sides by -1 \[x = 52\] That help?

OpenStudy (anonymous):

@Ryan88 in the end it's x=-52 not 52

OpenStudy (anonymous):

OK but @Abdulhameed said it was -52 and u ( @Ryan88 ) your saying its just 52 which is it?

OpenStudy (anonymous):

Oops, typo. @Abdulhameed is right. It's -52